You decide its time to start saving for your retirement by making equal bi-weekly deposits into an account that earns 8% year, compounded daily. How long will it take for you to have 100 times the amount of the initial deposit in your account?
Answer #1

Rate compounded biweekly=((1+8%/365)^(365/26)-1)*26=8.011%

x/(8.011%/26)*((1+8.011%/26)^t-1)=100*x

=>t=log(1+100*8.011%^26)/log(1+8.011%/26)

=>t=87.31

It will take 87.31/26=3.36 years